Out of 25 employees of a company, 5 are engineers. Three employees are selected at random for granting leave. What is the probability that (i) all the three are engineers?(ii) none of them is an  engineer? (iii) at least one of them is an engineer


1
Expert's answer
2021-05-03T07:15:12-0400

The total number of ways to select three employees out of 25 employees is


"\dbinom{25}{3}=\dfrac{25!}{3!(25-3)!}=2300"

(i)


"P(3\ engineers)=\dfrac{\dbinom{5}{3}\dbinom{25-5}{3-3}}{\dbinom{25}{3}}=\dfrac{10\cdot1}{2300}""=\dfrac{1}{230}"

(ii)



"P(0\ engineers)=\dfrac{\dbinom{5}{0}\dbinom{25-5}{3-0}}{\dbinom{25}{3}}=\dfrac{1\cdot1140}{2300}""=\dfrac{57}{115}"

(iii)


"P(at\ least \ 1\ engineer)=1-P(0\ engineers)""=1-\dfrac{57}{115}=\dfrac{58}{115}"